OpenStudy (anonymous):

if m and n are integers, find values of m and n such that m^2 and n^2 - 12

OpenStudy (anonymous):

... they what? .-.

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

if m and n are integers, find values of m and n such that m^2 and n^2 - 12 .... what should they equal, ... or what condition should they satisfy?

OpenStudy (anonymous):

they should both equal 21 when n(squared) is subtracted from m(squared)

OpenStudy (anonymous):

?

Parth (parthkohli):

They can't equal \(21\). They gotta be \(42\).

OpenStudy (anonymous):

so you want m^2 + n^2 = 21?

OpenStudy (anonymous):

how an example of such a solution is \[5^{2}-2^{2}=21\]

Parth (parthkohli):

So\[(m +n)(m - n) = 21\]

Parth (parthkohli):

Diophantine...

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

\(n+m=7,n-m=3\) solve for \(n,m\)

OpenStudy (anonymous):

since \(21=7\times 3\)

Parth (parthkohli):

Right there.

Parth (parthkohli):

\[n - m = 3, n + m = 7\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

also \(m+n=21,m-n=1\) will give another solution

Parth (parthkohli):

Oh I mean\[n + m = 3 , n - m = 3\]

OpenStudy (anonymous):

nope it is a quadratic expression

OpenStudy (anonymous):

i wish i knew what the original question actually said

OpenStudy (anonymous):

and it is m^2-n^2=21

OpenStudy (anonymous):

i just need numbers that make 21 when they are subtracted

Parth (parthkohli):

\[m^2 - n^2 = (m + n)(m - n)\]Proceed.

Parth (parthkohli):

Wait, numbers when they are subtracted?\[m -n = 21\]?

OpenStudy (anonymous):

the squares

OpenStudy (anonymous):

got it then what @ParthKohli said will work. solve the system (in your head) \[m+n=7\]\[m-n=3\]

Parth (parthkohli):

Eliminate, eliminate, eliminate, pirate!

OpenStudy (anonymous):

\(m=5,n=2\)

OpenStudy (anonymous):

that is one

OpenStudy (anonymous):

i need one more

OpenStudy (anonymous):

or \[m+n=21\]\[m-n=1\] \[m=11,m=10\]

Parth (parthkohli):

\[m - n = 21 \\ m +n = 1\]

OpenStudy (anonymous):

in other words there are two ways to factor \(21\) either use \(7\times 3\) or \(21\times 1\)

OpenStudy (anonymous):

@parthkohli i agree that it is 7 and 3 but i already got that i need one more example

OpenStudy (anonymous):

since \(m^2-n^2=(m+n)(m-n)\) you use either factorization

Parth (parthkohli):

Actually if you look,\[m = \pm 21, n = \pm 1\]

OpenStudy (anonymous):

wait i got it its \[\pm10\] and \[\pm11\]